in a class 65% students are boys and 35% students are girls the probability that a boy comes late for a lecture is 0.4 and the probability that a girls comes late for a lecture is 0.2. given that a student has come late what is the probability that the student is girl?
1
Expert's answer
2020-12-02T19:23:06-0500

Given that,


No of boys in a class = 65%

No of girls in a class = 35%


Probability of a boy coming late to the class = 0.4

Probability of a girl coming late to the class = 0.2


Hence the formula for probability for any event A given an event B = "\frac {P(A\bigcap B)}{P(B)}"


= "\frac {probability\ that \ a \ girl\ coming\ late}{probability\ that\ a\ student\ is\ late}"


="\frac {0.35*0.2}{0.65*0.4+0.35*0.2}=\frac {0.07}{0.33}=0.212121"